OpenStudy (anonymous):

Create a system of equations that includes one linear equation and one quadratic equation. y=x^2-4x-2 y=x-2 And after graphing and everything I got my solutions 0,-2 and 5,3 My teacher told me this Task two is good but you still need to make them equal to each other and solve, Can anyone help me with the final part.. Last question till i'm done! Thank you!!

OpenStudy (anonymous):

@Klip

OpenStudy (anonymous):

watch this it may help

OpenStudy (anonymous):

I already know what x and y intercepts are? She said I need to make them equal and solve, what do I need to make equal. @CrazyMexican816

OpenStudy (anonymous):

@CrazyMexican816 Could you help me?

OpenStudy (anonymous):

honestly im confused as well.. im sorry but maybe @Bigbosssaint21 could help you hes pretty smart.

OpenStudy (anonymous):

Okay thank you, i'll wait for him

OpenStudy (anonymous):

sorry i wasnt a help to you ...:( here have a medal

OpenStudy (anonymous):

you to

OpenStudy (anonymous):

thx I really appreciate it you have a good rest of the day

OpenStudy (anonymous):

@Bigbosssaint21 Could you help me

OpenStudy (anonymous):

@kaylamarie101 , @Acarpenter2 , @jcpd910 , @Jumperman , @ShailKumar plz help yuk15

OpenStudy (anonymous):

and @Law&Order

OpenStudy (anonymous):

@Law&Order

OpenStudy (anonymous):

y=x^2-4x-2 y=x-2 x^2 - 4x - 2 = 0 x - 2 = 0

OpenStudy (anonymous):

Now they are equal. :D

OpenStudy (anonymous):

@jcpd910 Thank you very much

OpenStudy (anonymous):

No problem.